# Forecasting Clinical Risk from Textual Time Series: Structuring Narratives for Temporal AI in Healthcare

### v3

@@ Discussion
Methodological Contributions and Potential Impact. Our framework demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks. The ability to extract temporally structured insights from unstructured clinical text could potentially support clinical decision-making, particularly in settings with limited access to structured data or specialized expertise.
@@
Our approach also demonstrates that our language model systems can reliably capture the temporal reasoning that clinicians use in practice based on their clinical documentation. However, additional work would be required to validate performance on real-time clinical data and demonstrate measurable impact on patient outcomes in deployment. Further discussion on societal impact of our work is in Appendix I.
Limitations and Future Directions. Our study has important limitations to consider. First, the pipeline relies on case reports from the PubMed Open Access (PMOA) corpus, which often reflect rare or atypical presentations and differ from routine clinical notes (e.g., progress notes, discharge summaries), potentially limiting the model’s generalizability to real-world health system texts. Compared to routine clinical notes, case reports offer rich, temporally structured narratives with explicit clinical rationale, making them especially useful for evaluating model performance under sparse documentation. This study develops a general framework for textual time series forecasting, using case reports as an interpretable, temporally rich foundation for method development with potential extensions to real-world clinical corpora like MIMIC discharge summaries.
@@
Third, despite efforts to mitigate causal leakage during training and evaluation (e.g., timestamp extraction, temporal masking, and use of external test sets), pretrained language models may have been exposed to PMOA content. Although these models were not trained on our derived (event, time) sequences or temporally reordered data, it is possible that the original case report narratives were included in their pretraining data. Our pipeline applies a series of transformations—including temporal reordering and masking of future events beyond a prediction cutoff—which fundamentally alter the input representation and enforce a causal structure that is not present in the original documents. This creates a distinct prediction setup that requires models to reason over partial, temporally grounded inputs rather than complete narratives. To further assess the impact of pretraining exposure, we include results from RedPajama-INCITE-7B-Instruct, an open-source model with a fully documented, PubMed-free training corpus. Its comparable performance to other decoder models (0.352 F1 at 24h) and similar concordance (0.618) to models potentially exposed to PubMed provides evidence that our T2S2 prediction task is sufficiently distinct from raw PubMed text, helping to mitigate causal leakage concerns. A detailed discussion on the scope of using open-source models for avoiding potential data leakage has been presented in Appendix H.
Finally, our analysis is limited by the modest size of the high-cost manual annotations collected as a gold standard to calibrate and validate our LLM extraction pipeline. Expanding gold-standard validation across diverse diagnoses could assess the generalizability of our approach and aid in evaluating model robustness, and is a direction for future work.
@@
  • Appendix A: Complete forecasting performance tables for all models on DeepSeek-R1 and Llama-3.3-70B annotations
  • Appendix B: Sensitivity analysis of masking history on Llama-3.3-70B annotations
  • Appendix C: Technical details of encoder-based forecasting methods, including MLP-head and masking approaches
  • Appendix D: Survival modeling framework using language model embeddings and the hyperparameters used
  • Appendix E: Computing infrastructure used for experiments, including hardware specifications, memory requirements, operating system, and software environment
@@
  • Appendix E: LLM prompt template for extracting clinical time series from case reports
  • Appendix E: Prompting strategies for zero-shot and few-shot forecasting
  • Appendix H: Analysis of open-source models to address data leakage concerns
  • Appendix I: Broader societal impact and ethical considerations
@@
Appendix A Forecasting performance on DeepSeek-R1 and Llama-3.3-70B textual time-series annotations
T2S2 (DeepSeek R1) sepsis-10 sepsis-100
@@
T2S2 (Llama 3.3) sepsis-10 sepsis-100
@@
Tables A1 and A2 present the forecasting performance of our models on textual time-series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively. We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
Appendix B Masking history: Llama-3.3-70B textual time-series annotations
Table A3 examines the impact of randomly masking portions of the input history on model performance when using Llama-3.3-70B-generated annotations. We evaluate performance across three datasets (T2S2-L33, Sepsis-10, and Sepsis-100) using F1 scores for prediction of next $k$ events happening with the next day and concordance index (c-index) correctly ordering the next $k$ events. As the proportion of masked history increases, we observe a general decline in F1 performance, particularly at higher masking rates (e.g., 90%), suggesting that access to historical context is important for accurate forecasting. Interestingly, the concordance index remains relatively stable, indicating the model’s robustness in ranking the next $k$ events with partial input histories.

### v5

@@ Results
Additional insights into the forecasting results are presented in Appendix J. Specifically, we summarize two observations that clarify model behavior and apparent anomalies: (i) why decoders fare better on survival but not short-horizon forecasting, and (ii) interpreting very low $F_{1}$ scores and external validation drops.
Methodological Contributions and Potential Impact. Our framework demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks. The ability to extract temporally structured insights from unstructured clinical text could potentially support clinical decision-making, particularly in settings with limited structured data or specialized expertise.
@@
Our approach also demonstrates that our language model systems can reliably capture the temporal reasoning that clinicians use in practice based on their clinical documentation. However, additional work would be required to validate performance on real-time clinical data and demonstrate measurable impact on patient outcomes in deployment. Further discussion on societal impact of our work is in Appendix M.
Limitations and Future Directions. Our study has important limitations to consider. First, our pipeline relies on case reports from the PubMed Open Access (PMOA) corpus. Because these reports often highlight rare or atypical presentations and differ from routine clinical notes (e.g., progress notes, discharge summaries), generalizability to real-world health-system text may be limited. However, case reports provide rich, temporally structured narratives with explicit clinical rationale, making them well suited for evaluating model performance under sparse documentation. Accordingly, we present a general framework for textual time series forecasting that uses case reports as an interpretable, temporally rich foundation for method development, with extensions to real-world corpora such as MIMIC discharge summaries. See Appendix L for a detailed discussion on the generalizability of our pipeline beyond case report details, as well as comparisons with other clinical sources.
@@
Finally, despite mitigation steps (timestamp extraction, temporal masking, external test sets), pretrained language models may still have been exposed to PMOA content. Although they were not trained on our derived (event, time) sequences or temporally reordered data, the original narratives could appear in pretraining. Our pipeline applies temporal reordering and masking of future events beyond the prediction cutoff, enforcing a causal structure absent from the originals and yielding a distinct setup that requires reasoning over partial, temporally grounded inputs. While we included RedPajama-INCITE-7B-Instruct, an open-source model with a documented PubMed-free corpus for pretraining, and saw its performance (0.352 F1 at 24h; concordance 0.618) was comparable to decoders potentially exposed to PubMed, other strategies supporting the argument that the T2S2 task is sufficiently distinct from raw PMOA text and mitigates causal leakage concerns could be examined. See Appendix K for a detailed discussion of how open-source models mitigate potential data leakage, as well as additional forecasting results on 115 case reports published after the pretraining cutoffs of encoder-based models (post-2020).
Acknowledgments
This research was supported in part by the Division of Intramural Research (DIR) of the National Library of Medicine (NLM), National Institutes of Health. This work utilized the computational resources of the NIH HPC Biowulf cluster. S.N. was supported by Carnegie Mellon University TCS Presidential Fellowship, and Natural Sciences and Engineering Research Council of Canada (NSERC) PGS-D award. S.N. was also supported in part by an appointment to the National Library of Medicine Research Participation Program administered by the Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S. Department of Energy (DOE) and the National Library of Medicine, National Institutes of Health. ORISE is managed by ORAU under DOE contract number DE-SC0014664. All opinions expressed in this paper are the authors’ and do not necessarily reflect the policies and views of NIH, NLM, DOE, or ORAU/ORISE.
@@ Acknowledgments
  • Appendix A: Validity of LLM-extracted temporal annotations (adopted from our prior work (adopted from our prior work (Wang and Weiss 2025)).
@@
  • Appendix A: LLM prompt template for extracting clinical time series from case reports.
  • Appendix A: Prompting strategies for zero-shot and few-shot forecasting.
  • Appendix D: Technical details of encoder-based forecasting methods, including MLP-head and masking approaches.
  • Appendix E: Survival modeling framework using language model embeddings and the hyperparameters used.
@@
  • Appendix E: Complete forecasting performance tables for all models on DeepSeek-R1 and Llama-3.3-70B annotations.
  • Appendix E: Complete survival analysis performance table for all datasets.
  • Appendix E: Sensitivity analysis of the forecasting performance when using time-ordered versus text-ordered training.
  • Appendix E: Sensitivity analysis of masking history on Llama-3.3-70B and DeepSeek-R1 annotations.
  • Appendix J: Additional insights into forecasting results.
  • Appendix K: Analysis of open-source models to address data leakage concerns.
  • Appendix L: Generalizability beyond PubMed case reports: comparison with other clinical sources.
  • Appendix M: Broader societal impact and ethical considerations.
Appendix A Validity of Temporal Annotations
We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports. We refer to this as Textual Time Series for Sepsis (T2S2). We assessed the quality of T2S2 using three key metrics: event match rate, temporal concordance (c-index), and log-time discrepancy on two external validation sets: sepsis-10 (n = 10) and sepsis-100 (n = 90).(adopted from our prior work (Wang and Weiss 2025)).
Text Event Match
To align predicted clinical findings with reference annotations, we use a recursive best-match procedure adapted from (Wang and Weiss 2025). At each step, we select the reference-prediction pair with the minimal string-based distance; ties are broken by text order (i.e., the order of events in the annotation files). Matched items are removed and the process recurses on the remainder, preventing many-to-one mappings. We evaluated Levenshtein distance, cosine similarity of BERT (bert-base-uncased) embeddings, and sentence-embedding cosine similarity from PubMedBERT (S-PubMedBert-MS-MARCO). Prior validation showed PubMedBERT-based cosine similarity with a threshold of 0.1 performed best, and we used this metric for evaluation. Pseudocode is provided in Algorithm 1.
Temporal Ordering
Temporal ordering accuracy was quantified with the concordance index (c-index), the probability that a randomly chosen pair of events is correctly ordered. For reference times $\{t^{r}_{i}\}$ and predicted times $\{t^{p}_{i}\}$,
$\text{c-index}=\frac{1}{N}\sum_{i<j}\mathds{1}\{(t^{r}_{i}-t^{r}_{j})(t^{p}_{i}-t^{p}_{j})>0\},$
where $N$ is the number of comparable pairs (event indices $i,j$ with $t^{r}_{i}\neq t^{r}_{j}$ and $t^{p}_{i}\neq t^{p}_{j}$), and $\mathds{1}\{\cdot\}$ is the indicator function. Higher values indicate better agreement with the reference timeline.
Algorithm 1 Recursive Best Match
1:Two lists: ref (reference events) and pred (predicted events)
2:List of best-matching event pairs
3:function MatchEvents(ref, pred)
4:  if ref is empty or pred is empty then
5:   return $[]$
6:  end if
7:  $\text{min\_distance}\leftarrow\infty$
8:  $\text{best\_pair}\leftarrow\text{None}$
9:  for all $r\in\texttt{ref}$ do
10:   for all $p\in\texttt{pred}$ do
11:     $d\leftarrow\text{ComputeDistance}(r,p)$
12:     if $d<\text{min\_distance}$ then
13:      $\text{min\_distance}\leftarrow d$
14:      $\text{best\_pair}\leftarrow(r,p)$
15:     else if $d=\text{min\_distance}$ then
16:      Get indices of $r$, $p$, and current best_pair
17:      if index of $r<$ best $r$ index then
18:        $\text{best\_pair}\leftarrow(r,p)$
19:      else if index of $r=$ best $r$ index and index of $p<$ best $p$ index then
20:        $\text{best\_pair}\leftarrow(r,p)$
21:      end if
22:     end if
23:   end for
24:  end for
25:  Remove best_pair.r from ref
26:  Remove best_pair.p from pred
27:  $\text{result}\leftarrow[\text{best\_pair}]+\text{MatchEvents}(\texttt{ref},\texttt{pred})$
28:  return result
29:end function
Time Discrepancy Analysis
Temporal prediction accuracy was assessed via discrepancies between predicted and reference timestamps. We computed the log-time discrepancy as $\log(1+\Delta t)$, where $\Delta t$ is the absolute time difference (in hours), and summarized errors with the Area Under the Log-Time Cumulative Distribution Function (AULTC), which reflects how rapidly the error CDF rises—larger AULTC indicates better temporal alignment.
sepsis-10 Model Event Match Rate Concordance AULTC DeepSeek-R1 0.79 0.876 0.772 Llama-3.3-70B 0.75 0.932 0.759 sepsis-100 Model Event Match Rate Concordance AULTC DeepSeek-R1 0.91 0.875 0.801 Llama-3.3-70B 0.78 0.894 0.791
Table A1: Comparison of event match rate (Event), median concordance (c), median absolute error (MAE), and area under the log-time curve (AULTC) across DeepSeek-R1 and Llama-3.3-70B for sepsis-10, and sepsis-100 datasets.
Formally, we compute the AULTC as the area under $F(x)$ from $x=0$ to $x=\log(1+S_{\text{max}})$, normalized by $\log(1+S_{\text{max}})$:
AULTC $=\frac{1}{\log(1+S_{\max})}\sum_{i=1}^{k}\big(x_{(i)}-x_{(i-1)}\big)\frac{i}{k}$
$\quad+\frac{\log(1+S_{\max})-x_{(k)}}{\log(1+S_{\max})},$
where $x_{(0)}=0$. With this definition, AULTC = 1 indicates that discrepancies are zero (perfect recovery), resulting in maximum area $\log(1+S_{\text{max}})$, and AULTC = 0 indicates that all discrepancies exceed $S_{\text{max}}$, yielding zero area. We also stratified results by reference time intervals (hour, day, week, year) to examine accuracy as a function of time from case presentation.
Annotation Quality Results
On the sepsis-10 subset, DeepSeek-R1 achieved an event match rate of 0.79, c-index of 0.876, and AULTC of 0.772, while Llama-3.3-70B obtained 0.75, 0.932, and 0.759, respectively. Thus, Llama provided stronger pairwise ordering (higher c-index) but with fewer matched events and a slightly weaker log-time error profile (lower AULTC) than DeepSeek. On sepsis-100, DeepSeek-R1 reached 0.91 (event match rate), 0.875 (c-index), and 0.801 (AULTC), whereas Llama-3.3-70B reached 0.78, 0.894, and 0.791. Here, DeepSeek matched substantially more events and achieved better AULTC, while Llama again showed a small c-index advantage. Overall, DeepSeek tends to optimize event identification and log-time alignment, whereas Llama slightly favors pairwise temporal ordering.
Overall, these metrics together indicate strong alignment of the generated textual time series with expert annotations, sufficient to support forecasting benchmarks.
B  LLM Prompt to Generate Clinical Textual Time Series from PMOA Case Reports
Prompt You are a physician. Extract the clinical events and the related time stamp from the case report. The admission event has timestamp 0. If the event is not available, we treat the event, e.g. current main clinical diagnosis or treatment with timestamp 0. The events that happened before event with 0 timestamp have negative time, the ones after the event with 0 timestamp have positive time. The timestamps are in hours. The unit will be omitted when output the result. If there is no temporal information of the event, please use your knowledge and events with temporal expression before and after the events to provide an approximation. We want to predict the future events given the events happened in history. For example, here is the case report. An 18-year-old male was admitted to the hospital with a 3-day history of fever and rash. Four weeks ago, he was diagnosed with acne and received the treatment with minocycline, 100 mg daily, for 3 weeks. With increased WBC count, eosinophilia, and systemic involvement, this patient was diagnosed with DRESS syndrome. The fever and rash persisted through admission, and diffuse erythematous or maculopapular eruption with pruritus was present. One day later the patient was discharged.
18 years old | 0
male | 0
admitted to the hospital | 0
fever | -72
rash | -72
acne | -672
minocycline | -672
increased WBC count | 0
eosinophilia | 0
systemic involvement | 0
diffuse erythematous or maculopapular eruption | 0
pruritus | 0
DRESS syndrome | 0
fever persisted | 0
rash persisted | 0
discharged | 24
C  Prompting Strategies for Forecasting Tasks
Zero-Shot
Prompt You are an expert physician. Reply to the prompt with structured predictions in a k-item, bar-separated row. For example, if there are k=3 events (A, B, C) and only B occurs in the time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response. Reply with the answer only. Do NOT provide any other text. Do NOT explain.
Zero-Shot (with degree of confidence)
Prompt You are an expert physician. Reply to the prompt with structured predictions in a k-item, bar-separated row. Assess whether the event happens in the time window using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen. For example, if there are k=3 events (A, B, C) and only B occurs in the time window, then: $1|5|1$ would be the correct response. Reply with the answer only. Do NOT provide any other text. Do NOT explain.
Few-Shot
Prompt You are an expert physician. Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window. Example 1. If there are k=3 events (A, B, C) and only B occurs in the requested time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response. Example 2. Suppose the time series is: admitted | 0
history of diabetes | 0
elevated troponin | 5
--- PREDICTION WINDOW ---
cardiac catheterization | 6
admission to icu | 24
transfer to floor | 72
discharge | 144
office visit | 1440
Few-Shot Ordinal
Prompt You are an expert physician. Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window, using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen. Example 1. If there are k=3 events (A, B, C) and only B occurs in the time window, then: 1 $\mid$ 5 $\mid$ 1 would be the correct response. Example 2. Suppose the time series is: admitted | 0
history of diabetes | 0
elevated troponin | 5
--- PREDICTION WINDOW ---
cardiac catheterization | 6
admission to icu | 24
transfer to floor | 72
discharge | 144
office visit | 1440
Appendix D Textual Time Series Forecasting Framework of Encoder Models
Input Preprocessing and Timeseries Construction.
Model Input Representation.
Language Models.
  • bert-base-uncased (BERT; Devlin et al., 2018)
  • roberta-base (Liu et al., 2019)
  • microsoft/deberta-v3-small (He et al., 2021)
  • answerdotai/ModernBERT-base, answerdotai/ModernBERT-large
  • thomas-sounack/BioClinical-ModernBERT-base, thomas-sounack/BioClinical-ModernBERT-large
Prediction Tasks.
1. Event Ordering (Concordance) Task.
  • Input: The serialized history segment followed by:
  • Label: $y=1$ if $\tau_{i}<\tau_{j}$, else $y=0$
2. Time-Window Classification Task.
  • Input: The serialized history segment followed by:
    $\texttt{[SEP]}\;\tau_{j}\texttt{:}f_{j}\;\texttt{[SEP]}$ (3)
Masked Language Modeling Approaches.
1. Masked Event Occurrence Prediction.
  • Target: The model predicts either ”yes” or ”no” at the [MASK] position
2. Masked Temporal Ordering.
  • Input: History followed by: [SEP] $e_{1}$ [MASK] $e_{2}$ [SEP]
Implementation Details.
Training and Evaluation.
Appendix E Survival Modeling with Language Model Embeddings
Table A2: Hyperparameter grid search ranges for our survival models.
F  Forecasting Performance on DeepSeek-R1 and Llama-3.3 Textual Time Series Annotations
Tables A3 and A4 present the forecasting performance of our models on textual time series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively. We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
T2S2 (DeepSeek-R1) sepsis-10 sepsis-100
@@ Appendix E Survival Modeling with Language Model Embeddings
T2S2 (Llama-3.3) sepsis-10 sepsis-100
@@
G  Survival Performance on DeepSeek-R1 Annotations for T2S2, sepsis-10, and sepsis-100
In this section, we report the full concordance analyses for the sepsis-10 and sepsis-100 annotation sets in addition to T2S2 test set, presented in Table A5. These results are the extended version of the T2S2-only table (Table 2) included in the main text.
Model T2S2 sepsis-10 sepsis-100
Time Threshold 0h 24h 168h 0h 24h 168h 0h 24h 168h
bert-base-uncased 0.60 0.61 0.53 1.00 1.00 1.00 0.58 0.62 0.58
roberta-base 0.55 0.60 0.60 0.75 1.00 1.00 0.55 0.58 0.57
deberta-v3-small 0.56 0.57 0.57 0.75 0.00 0.75 0.52 0.49 0.51
ModernBERT-base 0.52 0.58 0.58 1.00 1.00 1.00 0.54 0.51 0.51
ModernBERT-large 0.53 0.59 0.58 0.50 1.00 1.00 0.59 0.55 0.49
DeepSeek-R1-Llama-70B 0.64 0.63 0.59 0.50 1.00 1.00 0.61 0.53 0.55
Llama-3.3-70B-Instruct 0.62 0.63 0.58 0.50 1.00 0.75 0.64 0.57 0.53
DeepSeek-R1-Llama-8B 0.60 0.58 0.58 1.00 1.00 1.00 0.53 0.66 0.65
Llama-3.1-8B-Instruct 0.62 0.61 0.61 1.00 1.00 1.00 0.60 0.59 0.61
OLMO 32B Instruct 0.60 0.62 0.59 0.75 0.75 0.75 0.66 0.58 0.56
MediPhi-PubMed 0.54 0.62 0.55 1.00 1.00 1.00 0.62 0.58 0.61
Table A5: Survival analyses evaluated using time-dependent concordance index for each model across DeepSeek-R1 test set annotations (T2S2, sepsis-10, sepsis-100), evaluated at three observation time thresholds (0h, 24h, and 168h).
H  Sensitivity Analysis of Time-Ordered versus Text-Ordered Training
Table A6 shows time-ordering generally improves concordance, particularly for DeepSeek-R1 annotations, while F1 scores show mixed results. Text-ordering sometimes yields better external dataset performance with Llama-3.3-70B, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order. These results highlight a trade-off between temporal alignment and narrative coherence for different tasks.
F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
T2S2 (DeepSeek-R1)
Text-ordered 0.290 0.569 0.876 0.659 0.361 0.598 0.615 0.609
Time-ordered 0.306 0.576 0.877 0.672 0.428 0.576 0.607 0.613
T2S2 (Llama-3.3-70B)
Text-ordered 0.311 0.648 0.898 0.586 0.423 0.526 0.662 0.541
Time-ordered 0.257 0.645 0.903 0.632 0.382 0.595 0.616 0.612
Table A6: Comparison of text-ordered and time-ordered training for DeepSeek-R1 and Llama 3.3-70B annotations on F1 scores and c-index values.
I  Masking History: Llama-3.3-70B Textual Time Series Annotations
Tables A7 and A8 report the effect of timestep dropout on classification and ranking performance. We evaluate performance across three datasets (T2S2-L3.3, sepsis-10, and sepsis-100) using F1 scores for prediction of next $k$ events happening within the next day and concordance index (c-index) for correctly ordering the next $k$ events. As the proportion of masked history increases, we observe a general decline in F1 performance, particularly at higher masking rates (e.g., 90%), suggesting that access to historical context is important for accurate forecasting. Interestingly, the concordance index remains relatively stable, indicating the model’s robustness in ranking the next $k$ events with partial input histories.
More specifically, we can see that F1 scores decrease monotonically with increasing TDR, with relatively mild degradation up to 30–60% dropout and sharper declines thereafter. For example, in the DeepSeek-R1 annotations (Table A7), T2S2/sepsis-10/sepsis-100 drop from 0.576/0.428/0.607 at TDR = 0% to 0.423/0.200/0.510 at TDR = 90%, with sepsis-10 showing the largest relative decline. In contrast, concordance (c-index) remains stable across dropout levels (0.661–0.678), indicating that ranking-based evaluations are less sensitive to history truncation. These trends are consistent for Llama-3.3 annotations (Table A8), suggesting that ranking objectives capture temporal relationships that are more robust to missing historical information than binary classification accuracy of predicting whether an event happens within a time window.
Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
0% 0.576 0.672 0.428 0.576 0.607 0.613
15% 0.572 0.667 0.415 0.564 0.604 0.616
30% 0.548 0.672 0.357 0.542 0.591 0.619
60% 0.542 0.678 0.310 0.578 0.588 0.616
90% 0.423 0.661 0.200 0.561 0.510 0.614
Table A7: Effect of Randomly Masking History (DeepSeek-R1 annotations) on F1 and Concordance. Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively. c10/100: c-index for sepsis-10 and sepsis-100 respectively.